OpenStudy (anonymous):

Write the quadratic equation whose roots are 1 and -1 , and whose leading coefficient is 5

Directrix (directrix):

If the roots of the quadratic are 1 and -1, then that means that (x - 1) and (x + 1) are factors of the quadratic. That would render y = (x - 1)(x + 1). For the lead coefficient of 5, put a "5" out front as a factor. Y = 5*(x - 1)(x + 1). @darkarm146 Your task is to multiply out Y = 5*(x - 1)(x + 1) and get this quadratic in expanded form.